Circular No. 6297

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1996 B1

The following precise CCD positions have been reported:

1996 UT α_{2000} δ_{2000} m_{2} Observer
Jan. 29.30575 14^{h}01^{m}52.^{s}13 +53^{o}32′39′′.1 14.3 Offutt
29.34873 14 01 43.11 +53 32 11.0 14.2  _{”}
29.54324 14 01 02.86 +53 29 49.0  Kojima
29.55270 14 01 00.95 +53 29 42.8  _{”}
29.68130 14 00 33.90 +53 28 16.6  Kobayashi
29.68812 14 00 32.50 +53 28 12.5  _{”}
29.72067 14 00 25.62 +53 27 50.5  Asami
29.74031 14 00 21.10 +53 27 37.1  _{”}
29.78403 14 00 11.82 +53 27 07.5  Kushida
29.79757 14 00 08.81 +53 26 57.9  _{”}
29.87426 13 59 53.10 +53 25 52.1  Pravec
29.88146 13 59 51.63 +53 25 47.0  _{”}

W. Offutt (Cloudcroft). 0.6-m f/7 Ritchey-Chrétien. T. Kojima (Chiyoda). 0.25-m f/6.0 reflector. Strong condensation. J. Kobayashi (Kumamoto). 0.41-m f/5.0 reflector. Visually m_{1} = 8.8. A. Asami (Hadano). 0.20-m f/4.0 reflector. Y. and R. Kushida (Yatsugatake). 0.25-m f/6.3 Schmidt-Cassegrain. P. Pravec (Ondřejov). 0.65-m f/3.6 reflector.

EUVE J2115−58.6

N. Craig, University of California at Berkeley, reports: “An optical counterpart for EUVE J2115−58.6 has been identified as a magnetic CV, probably an AM Her-type cataclysmic variable. The field, centered at α = 21^{h}15^{m}41^{s}, δ = −58^{o}40′48′′ (equinox 2000.0), was observed as part of the EUVE Optical ID Campaign on 1995 Dec. 10.063 UT with the 1.5m telescope (Loral) at Cerro Tololo. The proposed optical counterpart, at α = 21^{h}15^{m}41^{s}, δ = −58^{o}40′54′′, well within the source error circle, is a star of estimated m_{v} = 16.0. The optical spectrum clearly shows the characteristics of a cataclysmic-variable spectrum, with the typical Balmer emission lines, as well as lines of He I, Ca II and strong He II emission. The continuum increases sharply toward the blue at short wavelengths. EUVE J2115−58.6 was detected during the all-sky survey in the EUVE Lexan/B band (5.8–17.4 nm, 0.07–0.21 keV) on 1992 Oct. 20, and also during the gap-filling on 1993 Apr. 19. A countrate of 0.05 cts/s is calculated. Monitoring photometric and spectroscopic observations are urged.”

1996 January 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden